A beam of X-rays is scattered by a target. At 45 degrees from the beam direction, the scattered x-rays have a wavelength of 2.2 pm. What is the wavelength of the x-rays in th direct beam

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Mahima
 
φ=45°λ’=2.2pmh/m0c=2.426×10-12λ’-λ=h(1-cosφ)/m0c
λ=1.54pm
 
Hope it helps
 
Regards
Arun (askIITians forum expert)